factor the quadratic expression completely, and find the roots of the expression.
135x^2 - 222x + 91

Answer by venugopalramana(3286)

135x2 - 222x + 91
135X^2-105X-117X+91=0
15X(9X-7)-13(9X-7)=0
(15x- 13)(9x-7)=0
15X-13=0......THAT IS .....15X=13....X=13/15
OR
9X-7=0........THAT IS ......9X=7.......X=7/9